# The Curse of Depth in Large Language Models

### v1

@@ Abstract
In this paper, we introduce the Curse of Depth, a concept that highlights, explains, and addresses the recent observation in modern Large Language Models (LLMs) where nearly half of the layers are less effective than expected. We first confirm the wide existence of this phenomenon across the most popular families of LLMs such as Llama, Mistral, DeepSeek, and Qwen. Our analysis, theoretically and empirically, identifies that the underlying reason for the ineffectiveness of deep layers in LLMs is the widespread usage of Pre-Layer Normalization (Pre-LN). While Pre-LN stabilizes the training of Transformer LLMs, its output variance exponentially grows with the model depth, which undesirably causes the derivative of the deep Transformer blocks to be an identity matrix, and therefore barely contributes to the training. To resolve this training pitfall, we propose LayerNorm Scaling, which scales the variance of output of the layer normalization inversely by the square root of its depth. This simple modification mitigates the output variance explosion of deeper Transformer layers, improving their contribution. Our experimental results, spanning model sizes from 130M to 1B, demonstrate that LayerNorm Scaling significantly enhances LLM pre-training performance compared to Pre-LN. Moreover, this improvement seamlessly carries over to supervised fine-tuning. All these gains can be attributed to the fact that LayerNorm Scaling enables deeper layers to contribute more effectively during training. Our code is available at LayerNorm-Scaling.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Layerwise output variance. This figure compares the output variance across various layers for different setups: (1) Pre-LN; (2) Pre-LN with Scaled Initialization; and (3) LayerNorm Scaling. The experiments are conducted on the LLaM-130M model trained for 10,000 steps. The proposed LayerNorm Scaling effectively controls the variance across layers.
@@
Recent studies reveal that the deeper layers (Transformer blocks) in modern LLMs tend to be less effective than the earlier ones (Yin et al., 2024; Gromov et al., 2024; Men et al., 2024). On the one hand, this interesting observation provides an effective indicator for LLM compression. For instance, we can compress deeper layers significantly more (Yin et al., 2024; Lu et al., 2024; Dumitru et al., 2024) to achieve high compression ratios. Even more aggressively, entire deep layers can be pruned completely without compromising performance for the sake of more affordable LLMs (Muralidharan et al., 2024; Siddiqui et al., 2024).
@@
On the other hand, having many layers ineffective is undesirable as modern LLMs are extremely resource-intensive to train, often requiring thousands of GPUs trained for multiple months, let alone the labor used for data curation and administration (Achiam et al., 2023; Touvron et al., 2023). Ideally, we want all layers in a model to be well-trained, with sufficient diversity in features from layer to layer, to maximize the utility of resources (Li et al., 2024b). The existence of ill-trained layers suggests that there must be something off with current LLM paradigms. Addressing such limitations is a pressing need for the community to avoid the waste of valuable resources, as new versions of LLMs are usually trained with their previous computing paradigm which results in ineffective layers.
@@
To seek the immediate attention of the community, we introduce the concept of the Curse of Depth (CoD) to systematically present the phenomenon of ineffective deep layers in various LLM families, to identify the underlying reason behind it, and to rectify it by proposing LayerNorm Scaling. We first state the Curse of Depth below.
@@
The Curse of Depth. The Curse of Depth refers to the observed phenomenon where deeper layers in modern large language models (LLMs) contribute significantly less to learning and representation compared to earlier layers. These deeper layers often exhibit remarkable robustness to pruning and perturbations, implying they fail to perform meaningful transformations. This behavior prevents these layers from effectively contributing to training and representation learning, resulting in resource inefficiency.
Empirical Evidence of CoD. The ineffectiveness of deep layers in LLMs has been previously reported. Yin et al. (2024) found that deeper layers of LLMs can tolerate significantly higher levels of pruning compared to shallower layers, achieving high sparsity. Similarly, Gromov et al. (2024) and Men et al. (2024) demonstrated that removing early layers causes a dramatic decline in model performance, whereas removing deep layers does not. Lad et al. (2024) showed that the middle and deep layers of GPT-2 and Pythia exhibit remarkable robustness to perturbations such as layer swapping and layer dropping. Recently, Li et al. (2024a) highlighted that early layers contain more outliers and are therefore more critical for fine-tuning. While these studies effectively highlight the limitations of deep layers in LLMs, they stop short of identifying the root cause of this issue or proposing viable solutions to address it.
To demonstrate that the Curse of Depths is prevalent across popular families of LLMs, we conduct layer pruning experiments on various models, including LLaMA2-7/13B, Mistral-7B, DeepSeek-7B, and Qwen-7B. We measure performance degradation on the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021) by pruning entire layers of each model, one at a time, and directly evaluating the resulting pruned models on MMLU without any fine-tuning in Figure 2.
Results: 1). Most LLMs utilizing Pre-LN exhibit remarkable robustness to the removal of deeper layers, whereas BERT with Post-LN shows the opposite trend. 2). The number of layers that can be pruned without significant performance degradation increases with model size.
@@
Identifying the Root Cause of CoD. We theoretically and empirically identify the root cause of CoD as the use of Pre-Layer Normalization (Pre-LN) (Baevski and Auli, 2019; Dai et al., 2019), which normalizes layer inputs before applying the main computations, such as attention or feedforward operations, rather than after. Specifically, while stabilizing training, we observe that the output variance of Pre-LN accumulates significantly with layer depth (see Appendix C), causing the derivatives of deep Pre-LN layers to approach an identity matrix. This behavior prevents these layers from introducing meaningful transformations, leading to diminished representation learning.
@@
Mitigating CoD through LayerNorm Scaling. We propose LayerNorm Scaling, which scales the output of Layer Normalization by the square root of the depth $\frac{1}{\sqrt{l}}$. LayerNorm Scaling effectively scales down the output variance across layers of Pre-LN, leading to considerably lower training loss and achieving the same loss as Pre-LN using only half tokens. Figure 1 compares the layerwise output variance across different setups: (1) Pre-LN, (2) Pre-LN with Scaled Initialization (Takase et al., 2023b), and (3) LayerNorm Scaling. As shown, Pre-LN exhibits significant variance explosion in deeper layers. In contrast, LayerNorm Scaling effectively reduces output variance across layers, enhancing the contribution of deeper layers during training. This adjustment leads to significantly lower training loss compared to Pre-LN. Unlike previous LayerNorm variants (Li et al., 2024b; Liu et al., 2020), LayerNorm Scaling is simple to implement, requires no hyperparameter tuning, and introduces no additional parameters during training. Furthermore, we further show that the model pre-trained with LayerNorm Scaling achieves better performance on downstream tasks in self-supervised fine-tuning, all thanks to the more effective deep layers learned.
Figure 2: Performance drop of layer pruning across different LLMs. (a) BERT-Large (Post-LN), (b) Mistral-7B (Pre-LN), (c) Qwen-7B (Pre-LN), (d) DeepSeek-7B (Pre-LN), (e) LLaMA2-7B (Pre-LN), and (f) LLaMA2-13B (Pre-LN). The results show that Pre-LN models exhibit significant inefficiency in deeper layers, while Post-LN models maintain strong deep-layer contributions.
Contributions.
  • We introduce the Curse of Depth to highlight, understand, and rectify the phenomenon in LLMs that is commonly overlooked—deep layers fail to contribute as effectively as they should.
  • We identify the root cause as Pre-LN, which causes output variance to grow exponentially with model depth. This leads to deep Transformer blocks having derivatives close to the identity matrix, rendering them ineffective during training. While scaled initialization (Shoeybi et al., 2020) helps mitigate variance at initialization, it does not prevent explosion during training.
  • To mitigate this issue, we propose LayerNorm Scaling, which inversely scales the output of Pre-LN by the square root of the depth. This adjustment ensures that all layers contribute effectively to learning, thereby improving LLM performance.
  • We hope this work brings greater attention to this issue, contributes to the improvement of LLMs, and maximizes the utilization of computational resources dedicated to training large models.
@@ 2 Empirical Evidence of the Curse of Depth
To empirically analyze the impact of layer normalization on the Curse of Depth in LLMs, we conduct a series of evaluations inspired by Li et al. (2024b), extending their methodology to compare Pre-LN and Post-LN models.
2.1 Experimental Setup
@@
Methods: We evaluate Pre-LN and Post-LN models by assessing the impact of layer pruning at different depths. Our hypothesis is that Pre-LN models exhibit diminishing effectiveness in deeper layers, whereas Post-LN has less effective early layers. To verify this, we empirically quantify the contribution of individual layers to overall model performance across a diverse set of LLMs.
@@
LLMs: We conduct experiments on multiple widely adopted LLMs: BERT-Large (Devlin, 2019), Mistral-7B (Jiang et al., 2023), LLaMA2-7B/13B (Touvron et al., 2023), DeepSeek-7B (Bi et al., 2024), and Qwen-7B (Bai et al., 2023). These models were chosen to ensure architectural and application diversity. BERT-Large represents a Post-LN model, whereas the rest are Pre-LN-based. This selection enables a comprehensive evaluation of the effects of layer normalization across varying architectures and model scales.
@@
Evaluation Metric: To empirically assess the impact of deeper layers in LLMs, we adopt the Performance Drop metric $\Delta P^{(\ell)}$, inspired by Li et al. (2024b). This metric quantifies the contribution of each layer by measuring the degradation in model performance following its removal. Specifically, it is defined as:
$\Delta P^{(\ell)}=P_{\text{pruned}}^{(\ell)}-P_{\text{original}},$ (1)
where $P_{\text{original}}$ represents the performance of the unpruned model, and $P_{\text{pruned}}^{(\ell)}$ denotes the performance after removing layer $\ell$. A lower $\Delta P^{(\ell)}$ suggests that the pruned layer plays a minor role in the model’s overall effectiveness. For BERT-Large, we evaluate performance on the SQuAD v1.1 dataset (Rajpurkar, 2016), which measures reading comprehension. For Mistral-7B, LLaMA2-13B, and Qwen-7B, we assess model performance on the MMLU benchmark (Hendrycks et al., 2021), a widely-used dataset for multi-task language understanding.
2.2 Layer Pruning Analysis
Figure 2 presents the performance drop ($\Delta P^{(\ell)}$) across different layers for six LLMs, including one Post-LN model (BERT-Large) and five Pre-LN models (Mistral-7B, LLaMA2-13B, Qwen-7B, DeepSeek-7B and LLaMA2-7B).
As shown in Figure 2 (a), pruning deeper layers in BERT-Large leads to a significant decline in accuracy on SQuAD v1.1, while pruning earlier layers has minimal impact. The performance drop $\Delta P^{(\ell)}$ becomes particularly severe beyond the 10th layer, highlighting the crucial role of deeper layers in maintaining overall performance in Post-LN models. In contrast, removing layers in the first half of the network results in negligible changes, indicating their limited contribution to the final output.
However, as shown in Figure 2 (b)-(f), Pre-LN models exhibit a contrast pattern, where deeper layers contribute significantly less to the overall model performance. For instance, as shown in Figure 2 (b) and (c), pruning layers in the last third of Mistral-7B and Qwen-7B results in a minimal performance drop on MMLU, indicating their limited contribution to overall accuracy. In contrast, pruning the first few layers leads to a substantial accuracy degradation, highlighting their crucial role in feature extraction. Similarly, Figure 2 (d) and (e) show that DeepSeek-7B and LLaMA2-7B follow a similar pattern, where deeper layers have little impact on performance, while earlier layers play a more significant role. Finally, as shown in Figure 2 (f), more than half of the layers in LLaMA2-13B can be safely removed. This suggests that as model size increases, the contrast between shallow and deep layers becomes more pronounced, with earlier layers playing a dominant role in representation learning. This observation underscores the need for the community to address the Curse of Depth to prevent resource waste.
@@ 3 Analysis of the Curse of Depth
3.1 Preliminaries
This paper primarily focuses on Pre-LN Transformer (Baevski and Auli, 2019; Dai et al., 2019). Let $x_{\ell}\in\mathbb{R}^{d}$ be the input vector at the $\ell$-th layer of Transformer, where $d$ denotes the feature dimension of each layer. For simplicity, we assume all layers to have the same dimension $d$. The layer output $y$ is calculated as follows:
@@ 3.2 Pre-LN Transformers
where the growth of $\sigma^{2}_{x_{l}}$ is sub-exponential, as shown by the following bounds:
@@
Based on Assumption 1 and the work of (Takase et al., 2023b), we obtain the following:
Theorem 1.
@@
The detailed description of $A$ and $B$, as well as the complete proof, are provided in Appendix A.2.
From Theorem 1, we observe that when the variance grows exponentially, as the number of layers $L\to\infty$, the norm $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ is bounded above by a fixed constant $M$. This result implies that even an infinitely deep Transformer remains stable, and by the Weierstrass Theorem, the network is guaranteed to converge. Consequently, this implies that for very large $L$, deeper layers behave nearly as an identity map from $x_{\ell}$ to $y_{\ell}$, thereby limiting the model’s expressivity and hindering its ability to learn meaningful transformations.
This outcome is undesirable, therefore, we would instead prefer the variance to increase more gradually—e.g., linearly—so that $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ exhibits linear growth. This observation highlights the necessity of appropriate variance control mechanisms, such as scaling strategies, to prevent excessive identity mappings and enhance network depth utilization.
3.3 Post-LN Transformers
For Post-LN Transformers, we continue to adopt Assumption 1. In this setting, each layer is followed by a layer normalization (LN) step, ensuring that the variances $\sigma^{2}_{x_{\ell}}$ and $\sigma^{2}_{x^{\prime}_{\ell}}$ remain fixed at 1 across all layers. Consequently, the norm $\left\|\frac{\partial y_{\ell}}{\partial x_{\ell}}\right\|_{2}$ exhibits minimal variation from one layer to the next, indicating stable gradient propagation.
Since the variance is effectively controlled by LN in Post-LN Transformers, an explicit variance‐based analysis becomes less critical. Nonetheless, there remain other important aspects to investigate in deeper Post-LN architectures, such as the evolution of feature mappings and the behavior of covariance kernels over deep layers. These directions will be pursued in future work.
4 LayerNorm Scaling
Our theoretical and empirical analyses indicate that Pre-LN amplifies output variance, leading to the Curse of Depth and reducing the effectiveness of deeper layers. To mitigate this issue, we propose LayerNorm Scaling, a simple yet effective normalization strategy. The core idea of LayerNorm Scaling is to control the exponential growth of output variance in Pre-LN by scaling the normalized outputs according to layer depth. Specifically, we apply a scaling factor inversely proportional to the square root of the layer index to scale down the output of LN layers, stabilizing gradient flow and enhancing the contribution of deeper Transformer layers during training. LayerNorm Scaling is illustrated in Figure 3.
Figure 3: Comparison between Pre-LN (a) and LayerNorm Scaling (b). LayerNorm Scaling applies a scaling factor inversely proportional to the square root of the layer index $l$, preventing excessive variance growth and stabilizing training dynamics across layers.

### v2

@@ Abstract
In this paper, we introduce the Curse of Depth, a concept that highlights, explains, and addresses the recent observation in modern Large Language Models (LLMs) where nearly half of the layers are less effective than expected. We first confirm the wide existence of this phenomenon across the most popular families of LLMs such as Llama, Mistral, DeepSeek, and Qwen. Our analysis, theoretically and empirically, identifies that the underlying reason for the ineffectiveness of deep layers in LLMs is the widespread usage of Pre-Layer Normalization (Pre-LN). While Pre-LN stabilizes the training of Transformer LLMs, its output variance exponentially grows with the model depth, which undesirably causes the derivative of the deep Transformer blocks to be an identity matrix, and therefore barely contributes to the training. To resolve this training pitfall, we propose LayerNorm Scaling (LNS), which scales the variance of output of the layer normalization inversely by the square root of its depth. This simple modification mitigates the output variance explosion of deeper Transformer layers, improving their contribution. Across a wide range of model sizes (130M to 7B), our experiments show that LNS consistently outperforms previous normalization and scaling techniques in enhancing LLM pre-training performance. Moreover, this improvement seamlessly carries over to supervised fine-tuning. All these gains can be attributed to the fact that LayerNorm Scaling enables deeper layers to contribute more effectively during training. Our code is available at LayerNorm-Scaling.
**footnotetext: Equal contribution.footnotetext: Correspondence to: Shiwei Liu, shiwei.liu@maths.ox.ac.uk.
Figure 1: Left: Schematic diagrams of (a) Pre-LN and (b) LayerNorm Scaling. LayerNorm Scaling applies a scaling factor inversely proportional to the square root of the layer index $\ell$, preventing excessive variance growth. Right: Language modeling loss of scaling up parameter count up to 7B. All models are trained for 20B tokens using OLMo (Groeneveld et al., 2024).
@@ 1 Introduction
Recent studies reveal that the deeper layers (Transformer blocks) in modern LLMs tend to be less effective than the earlier ones (Yin et al., 2024; Gromov et al., 2024; Men et al., 2024; Li et al., 2024b). On the one hand, this interesting observation provides an effective indicator for LLM compression. For instance, we can compress deeper layers significantly more (Yin et al., 2024; Lu et al., 2024; Dumitru et al., 2024) to achieve high compression ratios. Even more aggressively, entire deep layers can be pruned completely without compromising performance (Muralidharan et al., 2024; Siddiqui et al., 2024).
@@
On the other hand, having many layers ineffective is undesirable as modern LLMs are extremely resource-intensive to train, often requiring thousands of GPUs trained for multiple months, let alone the labor used for data curation and administration Achiam et al. (2023); Touvron et al. (2023). Ideally, we want all layers in a model to be well-trained, with sufficient diversity in features from layer to layer, to maximize the utility of resources (Li et al., 2024b). The existence of ill-trained layers suggests that there must be something off with current LLM paradigms. Addressing such limitations is a pressing need for the community to avoid the waste of valuable resources, as new versions of LLMs are usually trained with their previous computing paradigm which results in ineffective layers.
@@
To seek the immediate attention of the community, we re-introduce the concept of the Curse of Depth (CoD) to systematically present the phenomenon of ineffective deep layers in various LLM families, to identify the underlying reason behind it, and to rectify it by proposing LayerNorm Scaling. We first state the Curse of Depth below.
@@
The Curse of Depth. The Curse of Depth refers to the observed phenomenon where deeper layers in modern LLMs contribute significantly less (but not nothing) to learning and representation compared to earlier layers. These deeper layers often exhibit remarkable robustness to pruning and perturbations, implying they fail to perform meaningful transformations. This behavior prevents these layers from effectively contributing to training and representation learning, resulting in resource inefficiency.
Empirical Evidence of CoD. To demonstrate that CoD is a common phenomenon across prominent LLM families, we perform layer pruning experiments on Qwen3, LLaMA2, and DeepSeek. Specifically, we prune one layer at a time, without any fine-tuning, and directly evaluate the resulting pruned models on the MMLU benchmark (Hendrycks et al., 2021), as shown in Figure 2. Key findings: (1) Most models, including the latest Qwen3, exhibit surprising resilience to the removal of deeper layers; (2) The number of layers that can be removed without causing significant performance drop increases with model size; (3) Representations in deeper layers are significantly more similar to each other than those in earlier layers.
@@
Identifying the Root Cause of CoD. We theoretically and empirically identify the root cause of CoD as the use of Pre-Layer Normalization (Pre-LN) (Baevski and Auli, 2019; Dai et al., 2019), which normalizes layer inputs before applying the main computations, such as attention or feedforward operations, rather than after. Specifically, while stabilizing training, we observe that the output variance of Pre-LN accumulates significantly with layer depth as shown in Figure 4, causing the derivatives of deep Pre-LN layers to approach an identity matrix. This behavior prevents these layers from introducing meaningful transformations, leading to diminished representation learning.
@@
Mitigating CoD through LayerNorm Scaling. We propose LayerNorm Scaling (LNS), which scales the output of Layer Normalization by the square root of the depth $\frac{1}{\sqrt{l}}$. LayerNorm Scaling effectively scales down the output variance across layers of Pre-LN. LNS consistently delivers better pre-training performance than existing normalization and scaling techniques across various model sizes from 130M to 7B. Unlike previous LayerNorm variants (Li et al., 2024b; Liu et al., 2020), LayerNorm Scaling is simple to implement, requires no hyperparameter tuning, and introduces no additional parameters during training. [^1] We found that combining LNS with Scaled Initialization (Groeneveld et al., 2024; Radford et al., 2019; Shoeybi et al., 2020) diminishes the effectiveness of LNS. Therefore, we recommend removing the latter when applying LNS. Furthermore, we show that the model pre-trained with LayerNorm Scaling achieves better performance on downstream tasks in self-supervised fine-tuning, all thanks to the more diverse feature representations learned in deep layers.
Figure 2: Results of open-weight large-scale LLMs. Top: Performance drop after removing a single layer without fine-tuning. Bottom: Angular distance from the initial layer $\ell$ (x-axis) and its subsequent $n^{\text{th}}$ layer (y-axis). The results demonstrate that in Pre-LN LLMs, deeper layers produce highly similar representations to their adjacent layers, and their removal results in minimal performance degradation. In contrast, Post-LN models show the opposite trend: deep layers contribute more substantially to model performance.
Figure 3: Results of in-house small-scale LLaMa-130M. Angular Distance (a, b): Each column represents the angular distance from the initial layer $\ell$ (x-axis) and its subsequent $n^{th}$ layer (y-axis). The distance is scaled to the range [0, 1], where yellow indicates smaller distances and purple indicates larger distances. Performance Drop (c, d): ARC-e performance drop of removing each single layer from LLaMa-130M.
@@ 2 Empirical Evidence of the Curse of Depth
To empirically analyze the impact of layer normalization on the Curse of Depth in LLMs, we conduct a series of evaluations inspired by Li et al. (2024b), to compare Pre-LN and Post-LN models.
@@
Methodology: We evaluate Pre-LN and Post-LN models by assessing the impact of layer pruning at different depths. Our hypothesis is that Pre-LN models exhibit diminishing effectiveness in deeper layers, whereas Post-LN models have less effective early layers.
2.1 Open-weight Large-scale LLMs
@@
Models: To verify this, we empirically quantify the contribution of individual layers to overall model performance across a diverse set of LLMs, including Qwen3 (Team, 2025), LLaMA2 (Touvron et al., 2023), DeepSeek (Bi et al., 2024), and BERT-Large (Devlin, 2019). These models were chosen to ensure architectural and application diversity. BERT-Large represents a Post-LN model, whereas the rest are Pre-LN-based. This selection enables a comprehensive evaluation of the effects of layer normalization across varying architectures and model scales.
@@
Evaluation Metric: To empirically assess the impact of deeper layers in LLMs, we adopt two metrics, Performance Drop and Angular Distance, inspired by Gromov et al. (2024); Li et al. (2024b).
Performance Drop $\Delta P^{(\ell)}$ quantifies the importance of each layer by measuring the performance change after its removal. A smaller $\Delta P^{(\ell)}$ indicates that the pruned layer contributes less to the model’s overall performance. For BERT-Large, we evaluate using the SQuAD v1.1 dataset (Rajpurkar, 2016), while for other models, we use MMLU (Hendrycks et al., 2021), a standard benchmark for multi-task language understanding.
Angular Distance $d(x^{\ell},x^{\ell+n})$ quantifies the directional change between the input representations at layer $\ell$ and layer $\ell+n$ on a neutral pre-training dataset. Formally, given a token $T$, let $x_{T}^{\ell}$ and $x_{T}^{\ell+n}$ denote its input to layers $\ell$ and $\ell+n$, respectively. The angular distance is defined as:
$d(x^{\ell},x^{\ell+n})=\frac{1}{\pi}\arccos\left(\frac{x_{T}^{\ell}\cdot x_{T}^{\ell+n}}{\|x_{T}^{\ell}\|_{2}\|x_{T}^{\ell+n}\|_{2}}\right),$ (1)
where $\|\cdot\|_{2}$ denotes the $L^{2}$-norm. To reduce variance, we report the average distance over 256K tokens sampled from the C4 dataset. Smaller values of $d(x^{\ell},x^{\ell+n})$ indicate higher similarity between the two representations, suggesting limited transformation. Such layers can be considered redundant, as their removal minimally impacts the model’s internal representations. Ideally, each layer should introduce meaningful representational shifts to fully leverage the model’s capacity (Yang et al., 2023; Gromov et al., 2024).
Experimental Results: (1) Pruning deep layers in Pre-LN LLMs leads to negligible, and sometimes even positive, changes in performance, as shown in Figure 2-Top. Specifically, Figure 2 (b)–(d) reveals that a wide range of deeper layers—particularly beyond the 18th—can be pruned with minimal impact on performance. This indicates that deep layers in Pre-LN architectures contribute little to the model’s overall effectiveness. In contrast, Figure 2 (a) shows that pruning deep layers in BERT-Large (a Post-LN model) leads to a substantial drop in accuracy, while pruning early layers has a relatively minor effect. (2) Pre-LN models exhibit decreasing angular distance in deeper layers, indicating highly similar representations, as shown in Figure 2-Bottom. For instance, the angular distance in DeepSeek-7B falls below 0.2 after the 18th layer. Qwen3-8B demonstrates a higher similarity, with nearly half of its layers exhibiting distances below 0.2 from their preceding layers. In LLaMA2-13B, the angular distance approaches zero across the final one-third of the network. These similar representations align with the pruning results in Figures 2 (b)–(d), where pruning later layers has little effect, while pruning early layers significantly degrades performance.
Figure 4: Layerwise output variance. This figure compares the output variance across various layers for different setups: (1) Pre-LN; (2) Pre-LN with Scaled Initialization (Shoeybi et al., 2020; Radford et al., 2019); and (3) LayerNorm Scaling. The experiments are conducted on the LLaM-130M model trained for 10,000 steps. The proposed LayerNorm Scaling effectively controls the variance across layers.
2.2 In-house Small-scale LLaMa-130M
To eliminate the influence of other confounding variables, we train two LLaMA-130M models from scratch that differ only in their Layer Normalization, thereby clearly distinguishing Post-LN from Pre-LN, following Li et al. (2024b). The results are illustrated in Figure 3.
In Post-LN models, early layers exhibit high similarity (low angular distance, Figure 3-a) and their removal causes minimal performance loss (Figure 3-c), while deeper layers become more distinct and critical. Conversely, Pre-LN LLaMa-130M demonstrates a gradual decrease in angular distance with depth, resulting in highly similar deep layers (Figure 3-b). Removing most layers after the first in Pre-LN causes negligible performance loss (Figure 3-d), indicating their limited contribution. These consistent findings, observed in both open-weight and in-house LLMs, lead to the conclusion that the widespread use of Pre-LN is the root cause of the ineffectiveness of deep layers in LLMs.
@@ 3 Analysis of the Curse of Depth
Preliminaries. This paper primarily focuses on Pre-LN Transformer (Baevski and Auli, 2019; Dai et al., 2019). Let $x_{\ell}\in\mathbb{R}^{d}$ be the input vector at the $\ell$-th layer of Transformer, where $d$ denotes the feature dimension of each layer. For simplicity, we assume all layers to have the same dimension $d$. The layer output $y$ is calculated as follows:
@@ 3.1 Pre-LN Transformers
where the growth of $\sigma^{2}_{x_{\ell}}$ is sub-exponential, as shown by the following bounds:
@@
Based on Assumption 3.1 and the work of Takase et al. (2023b), we obtain the following:
Theorem 3.3.
@@
The detailed description of $A$ and $B$, as well as the complete proof, are provided in Appendix A.2. From Theorem 3.3, we observe that when the variance grows exponentially, as the number of layers $L\to\infty$, the norm $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ is bounded above by a fixed constant $M$. This result implies that even an infinitely deep Transformer remains stable, and by the Weierstrass Theorem, the network is guaranteed to converge. Consequently, this implies that for very large $L$, deeper layers behave nearly as an identity map from $x_{\ell}$ to $y_{\ell}$, thereby limiting the model’s expressivity and hindering its ability to learn meaningful transformations. This outcome is undesirable, therefore, we would instead prefer the variance to increase more gradually—e.g., linearly—so that $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ exhibits linear growth. This observation highlights the necessity of appropriate variance control mechanisms, such as scaling strategies, to prevent excessive identity mappings and enhance network depth utilization.
4 LayerNorm Scaling (LNS)
To mitigate the abovementioned issue, we propose LayerNorm Scaling, a simple yet effective normalization strategy. The core idea of LayerNorm Scaling is to control the exponential growth of output variance in Pre-LN by scaling the normalized outputs according to layer depth. Specifically, we apply a scaling factor inversely proportional to the square root of the layer index to scale down the output of LN layers, enhancing the contribution of deeper Transformer layers during training. LayerNorm Scaling is illustrated in Figure 1.